OpenStudy (anonymous):

hich of the following equations could be the result of using the comparison method to solve the system shown? x + y = 5 2x + y = 7 -x - 5 = 7 - 2x 5 - x = 7 - 2x 5 - x = 2x - 7

Parth (parthkohli):

1) subtract \(y\) from both equations. What do you get?

Parth (parthkohli):

No wait. Subtract \(x\).

Parth (parthkohli):

And then use substitution. Can you manage?

OpenStudy (anonymous):

i will work it out if i need more i will ask

OpenStudy (anonymous):

i got y=5-x 2x+5-x=7 so wat is the next step i mess up on this part

OpenStudy (anonymous):

-2x for 2x+5-x= 5 - x = 7 - 2x